# Dynamic neurons: A statistical physics approach for analyzing deep neural networks

### v1

@@ Abstract
Deep neural network architectures often consist of repetitive structural elements. We introduce a new approach that reveals these patterns and can be broadly applied to the study of deep learning. Similar to how a power strip helps untangle and organize complex cable connections, this approach treats neurons as additional degrees of freedom in interactions, simplifying the structure and enhancing the intuitive understanding of interactions within deep neural networks. Furthermore, it reveals the translational symmetry of deep neural networks, which simplifies the application of the renormalization group transformation—a method that effectively analyzes the scaling behavior of the system. By utilizing translational symmetry and renormalization group transformations, we can analyze critical phenomena. This approach may open new avenues for studying deep neural networks using statistical physics.
@@ I Introduction
In this paper, we develop a method to reformulate the cost function into a form that reveals symmetry, thereby reducing its complexity. We call this the ‘dynamic neuron’ approach. It can be understood as a decoupling technique that introduces additional degrees of freedom—the neurons in the bulk layers—coupled with the original degrees of freedom. After decoupling these degrees of freedom, the system exhibits translation symmetry similar to that of the Ising model. In a specific case, we show that this approach yields a Hamiltonian resembling that of the Ising model or the spin glass model. Due to this symmetry, many tools of statistical physics, including RG transformations and correlation lengths, become applicable to DNNs.
The remainder of this paper is organized as follows. In Sec. II, we provide a brief review of the deep learning and the statistical physics approach. In Sec III, we introduce the dynamic neuron approach and a new Hamiltonian to analyze DNNs. In Sec. IV, we present an example of the RG transformation using our new method. Critical phenomena are discussed in Sec. V. The summary and discussion are provided in Sec. VI, and a discussion of the spin glass model limit is included in App. A.
@@ II Fundamentals of Deep Learning
Figure 1: A deep neural network with $N=2$ and $M=2$, where $N$ is the width of each layer, and $M$ represents the depth of the DNN. For simplicity, the bias term $b$ has been omitted.
@@
A DNN is a system inspired by the brain [46, 47]. It consists of neurons, synaptic weights and synaptic biases, with neurons organized into layers. The $i$-th neuron in the $(m+1)$-th layer, denoted as $h_{i}^{(m+1)}$, is influenced by all neurons in the $m$-th layer, which are connected via synaptic weights $W_{ij}^{(m)}$ and synaptic biases $b_{i}^{(m)}$. (See Fig. 1.) This relationship is described by the following equation:
@@
In this formula, $\sigma$ represents a nonlinear function known as the activation function. [^1] Typically, the sigmoid function or ReLU is used as an activation function. An example of the sigmoid function is $\sigma(x)=\frac{1}{1+e^{-x}}$, and for ReLU, it is $\sigma(x)=\max(x,0)$. The repetitive calculation of Eq. (1) determines the value of each neuron, starting from the input (initial layer) $x=h^{(0)}$ and eventually reaching the output (final layer) $z=h^{(M)}$.
Simply put, deep learning is a method for approximating an unknown function using a DNN. The process of building a neural network from a given dataset is called training. There are several types of deep learning classified by their training methods, but we will focus on supervised learning for the sake of simplicity. In supervised learning, the machine is trained to give a good approximation for a set of examples, known as a training dataset. The training dataset $X$ consists of $L$ data vectors $x^{[l]}$, $l=1,2,\cdots,L$, and their corresponding label vectors $y^{[l]}$. For simplicity, we will often omit $[l]$, provided it does not cause confusion.
The goal of the training process is to minimize the discrepancy between the label $y^{[l]}$ and the output of the DNN $z^{[l]}$ by adjusting the synaptic weights $W_{ij}^{(m)}$, and biases $b_{i}^{(m)}$.
@@
A popular method for reducing the cost function is gradient descent. In each step of adjusting the synaptic weights and biases, the gradient of the cost function is calculated, and the synaptic weights and biases are adjusted by subtracting this value.
@@
From a physical perspective, Eq. (3) represents the equation of motion for the system [^3] In Eq. (3), the step size can be interpreted as the time interval between weight updates. Since it does not influence the underlying physics, we set $\eta=1$.. It is widely recognized that this equation of motion causes the system to conform to a Gibbs distribution [8]. Specifically, the probability distribution of encountering a particular synaptic configuration $W$, $b$ after a sufficiently long period of training is given by:
@@
where $\beta=\frac{1}{T}$, and $T$ is the temperature of the deep learning system, which depends on the noise in deep learning [8]. We incorporate the factor $\frac{1}{L}$ from $C$ into $\beta$. The partition function, denoted by $Z$, is defined as:
@@
where $\{W,b\}$ represents all possible configurations of synaptic weights and biases.
@@ III Dynamic neurons
To address this difficulty, we promote the neurons $h^{(m)[l]}$ to additional dynamic degrees of freedom and ensure that Eq. (1) is dynamically satisfied. Instead of using iterated activation functions, we introduce a series of Dirac delta functions and simplify the exponent to the absolute square as follows:
$\begin{split}Z=\int dWdbdh\prod_{i,l,m}\exp\Big{[}-&\beta\left(y^{[l]}-h^{(M)[l]}\right)^{2}\Big{]}\\ &\delta\left(h_{i}^{(m+1)[l]}-\sigma_{i}^{(m)[l]}\right).\end{split}$ (10)
Here, $i$ runs for the width, $l$ runs for the data number, and $m$ runs for the layer. We refer to these new dynamic degrees of freedom, $h^{(m)[l]}$, as “dynamic neurons.” While this partition function appears straightforward, it does not yet simplify the computation.
@@
$\begin{split}Z=\int d&Wdbdh\lim_{v\to 0^{+}}\left(\frac{1}{v\sqrt{2\pi}}\right)^{n}\\ &\exp\Big{[}-\beta\sum_{i,l}\left(y_{i}^{[l]}-\sigma^{(M)[l]}_{j}\right)^{2}\\ &\qquad\quad-\beta_{v}\sum_{i,l,m}\left(h^{(m+1)[l]}_{i}-\sigma_{i}^{(m)[l]}\right)^{2}\Big{]},\end{split}$ (12)
@@
We observe that incorporating the constraint terms reveals the internal structure of a DNN within the Hamiltonian framework. By defining the new Hamiltonian as
@@
with $\lambda_{m}=\beta_{v}/\beta$ except for $\lambda_{M}=1$. We can denote Eq. (12) in the typical way of describing the partition function,
@@
where the limit over $v$ is omitted for simplicity. The form of this new Hamiltonian is the summation of interactions between the nearest layers ($m$-th layer and $(m+1)$-th layer). In other words, interactions in the new Hamiltonian are restricted to the nearest degrees of freedom following the same pattern, unlike the original Hamiltonian.
@@
The emergence of this type of translation symmetry can simplify calculations and lead to interesting consequences. For example, in the one-dimensional and two-dimensional Ising models, the transfer matrix can be calculated explicitly due to translational symmetry. Therefore, it is expected that the new Hamiltonian can be analyzed by utilizing this symmetry. As an example, we perform the renormalization group transformation in Sec. IV.
@@
Moreover, this system can be effectively described using principles from statistical physics, which are often employed to uncover the macroscopic behavior of complex systems composed of similarly interacting units. With certain simplifying assumptions, this approach yields a Hamiltonian that closely resembles the spin glass model, a well-established framework in statistical physics. [^6] Several approaches have already introduced spin glass models to analyze DNNs under various assumptions [11, 12, 10, 13, 14, 15, 16, 17]. (See Appendix A). This approach offers a new method for applying the spin glass model to gradient descent-based training in DNNs.
@@ IV Renormalization Group Transformation
Since the new Hamiltonian resembles well-known statistical models, we can apply statistical physics tools to analyze the training of DNNs. A promising method in statistical physics is the renormalization group (RG) transformation, which is widely used in fields such as condensed matter and high-energy physics. This technique systematically simplifies a system by progressively integrating out smaller-scale fluctuations, leading to a renormalized system with fewer degrees of freedom.
@@
Figure 3: RG transformation applied to the DNN using the dynamic neuron approach, resulting in reduced network depth.
(i) The system consists of one neuron per layer ($N=1$) with a large depth ($M$ is large), and the training dataset contains a single data point ($L=1$). This assumption significantly reduces the system’s computational complexity. However, the following discussions can be easily extended to systems with larger $N$ or $L$, aside from the resulting increase in computational complexity.
@@
(iii) The weights $W$ and biases $b$ are restricted to values of $-1$ or $1$, and the neurons $h$ are limited to values of 00 or $1$. While, in theory, $h$ should be integrated over all possible values, we expect the Gaussian-like behavior of the partition function to constrain $h$ to values near 00 or $1$. Although in general deep neural networks, $W$ and $b$ can take arbitrary real numbers, we restrict their values here for simplicity. Interestingly, recent research has explored neural networks where parameters are similarly restricted to discrete values, such as $-1,1$ or $-1,0,1$. Despite these limitations, that model appears to achieve good performance [55, 56].
@@
As discussed in Sec. III, the Hamiltonian exhibits translational symmetry similar to that of the Ising model. Consequently, we can apply the RG transformation in a manner similar to its application in the Ising model. One approach involves summing over $W^{(m)}$, $b^{(m)}$, and $h^{(m)}$ for even-numbered layers in the partition function. After performing this summation, the partition function describes a system with fewer degrees of freedom. This real-space RG transformation method is known as decimation [57, 58], and the translational invariance plays a crucial role in performing the decimation. [^8] Decimation is generally avoided for systems with dimensions higher than one because it fails to account for the renormalization of the magnitude of the degrees of freedom, preventing it from achieving a true fixed point in higher dimensions.
@@
Figure 4: RG flow near the fixed points $e^{A}=1$, demonstrating that RG flows converge to the plane composed of fixed points.
@@
Let us discuss the physical meaning of the RG transformation in the dynamic neuron approach. As discussed earlier, the introduction of dynamic neurons results in additional terms, $\lambda_{m}H_{m}$, which reflect the constraints,
@@
can be interpreted as new constraints between the nearest degrees of freedom in the renormalized system. As the depth of the DNN decreases by applying the RG transformation, new constraints for the system with reduced depth emerge. Therefore, the RG transformation suggests statistically equivalent constraints for the system with reduced depth. Theoretically, by applying these new constraints to the partition function, the introduced dynamic neurons can be integrated out. Subsequently, a renormalized DNN system without dynamic neurons can be obtained. In general, after applying the RG transformation, the constraints often become more complex, leading to multiple solutions or analytically unsolvable equations. However, even if we cannot solve them explicitly, identifying these constraints can aid in understanding the structure of DNNs. In this sense, the dynamic neuron approach is a valuable tool for comprehending the nature of DNNs. This argument can be generalized without relying on the simplifying assumptions (i)–(iii).
The discussions in this section are schematically summarized in Fig. 6. Since the RG transformation is a useful technique in statistical physics, we aim to use it to analyze DNNs. However, the direct application of the RG transformation to DNNs is challenging, so we applied the dynamic neuron approach. Due to the translational symmetry that emerges from the dynamic neuron approach, the RG transformation can be applied much more easily. The RG transformation of a DNN using the dynamic neuron approach suggests new constraints for the system with reduced depth. By applying these constraints, a renormalized DNN system can be obtained.
DNNDNN with DyNRenormalizedDNNRenormalizedDNN with DyNRGRGIntroduceDyNApplyConstraintsRG
Figure 6: This diagram illustrates how to develop a renormalized DNN using the dynamic neuron (DyN) approach. Although not a rigorous explanation, it effectively conveys the essential concept. By introducing dynamic neurons, we obtain a new description of the DNN, and from the advantages of this new description, we derive a renormalized DNN with dynamic neurons. By applying constraints, the dynamic neurons can be integrated out, resulting in a renormalized DNN.

### v2

# Dynamic neuron approach to deep neural networks: Decoupling neurons for renormalization group analysis

@@ Abstract
Deep neural network architectures often consist of repetitive structural elements. We introduce an approach that reveals these patterns and can be broadly applied to the study of deep learning. Similarly to how a power strip helps untangle and organize complex cable connections, this approach treats neurons as additional degrees of freedom in interactions, simplifying the structure and enhancing the intuitive understanding of interactions within deep neural networks. Furthermore, it reveals the translational symmetry of deep neural networks, which simplifies the application of the renormalization group transformation—a method that effectively analyzes the scaling behavior of the system. By utilizing translational symmetry and renormalization group transformations, we can analyze critical phenomena. This approach may open new avenues for studying deep neural networks using statistical physics.
@@ I Introduction
In this paper, we develop a method to reformulate the cost function in a way that reveals its symmetry, thereby reducing its complexity. We refer to this as the ‘dynamic neuron approach’ (DNA). It can be viewed as a decoupling technique that introduces additional degrees of freedom—namely, neurons in the bulk layers—coupled to the original degrees of freedom. Once these degrees of freedom have been decoupled, the system exhibits translational symmetry akin to that of the Ising model. In a specific case, we show that this approach yields a Hamiltonian resembling that of either the Ising model or the spin-glass model. Due to this symmetry, many tools from statistical physics, including RG transformations and correlation lengths, become applicable to DNNs.
The remainder of this paper is organized as follows. In Sec. II, we provide a brief review of the deep learning and the statistical physics approach. In Sec III, we introduce the DNA and a new Hamiltonian to analyze DNNs. In Sec. IV, we present an example of the RG transformation using our new method. Critical phenomena are discussed in Sec. V. The summary and discussion are provided in Sec. VI, a discussion of the spin glass model limit is included in Appendix B, and the methods for implementing the dynamic neuron system are outlined in Appendix A.
@@ II Fundamentals of Deep Learning
Figure 1: A DNN with $N=2$ and $M=2$, where $N$ is the width of each layer, and $M$ represents the depth of the DNN. For simplicity, the bias term $b$ has been omitted.
@@
A DNN is a system inspired by the brain [46, 47]. It consists of neurons, synaptic weights, and synaptic biases, with neurons organized into layers. The $i$th neuron in the $(m+1)$th layer, denoted as $h_{i}^{(m+1)}$, is influenced by all neurons in the $m$th layer, which are connected via synaptic weights $W_{ij}^{(m)}$ and synaptic biases $b_{i}^{(m)}$. (See Fig. 1.) This relationship is described by the following equation:
@@
In this formula, $\sigma$ represents a nonlinear function known as the activation function. [^1] Typically, the sigmoid function or rectified linear unit(ReLU) is used as an activation function. An example of the sigmoid function is $\sigma(x)=\frac{1}{1+e^{-x}}$, and for ReLU it is $\sigma(x)=\max(x,0)$. The repetitive calculation of Eq. (1) determines the value of each neuron, starting from the input (initial layer) $x=h^{(0)}$ and eventually reaching the output (final layer) $z=h^{(M)}$.
Simply put, deep learning is a method for approximating an unknown function using a DNN. The process of building a neural network from a given dataset is called training. There are several types of deep learning classified by their training methods, but we will focus on supervised learning for the sake of simplicity. In supervised learning, the machine is trained to give a good approximation for a set of examples, known as a training dataset. The training dataset consists of $L$ data vectors $x^{[l]}$, $l=1,2,\cdots,L$, and their corresponding label vectors $y^{[l]}$. For simplicity, we will often omit $[l]$, provided it does not cause confusion.
The goal of the training process is to minimize the discrepancy between the label $y^{[l]}$ and the output of the DNN $z^{[l]}$ by adjusting the synaptic weights $W_{ij}^{(m)}$ and biases $b_{i}^{(m)}$.
@@
A popular method for reducing the cost function is gradient descent. In each step of adjusting the synaptic weights and biases, the gradient of the cost function is calculated, and the synaptic weights and biases are adjusted by subtracting this value:
@@
From a physical perspective, Eq. (3) represents the equation of motion for the system. [^3] In Eq. (3), the step size can be interpreted as the time interval between weight updates. Since it does not influence the underlying physics, we set $\eta=1$. To better capture the stochastic nature of training in deep learning, one may consider a Langevin-type equation of motion [48, 49, 50, 51, 52], where the update of weights includes both deterministic and stochastic components:
$\frac{dw}{dt}=-\nabla_{w}C+\eta(t),$ (4)
where $\eta(t)$ represents a white Gaussian noise with zero mean and time correlation $\langle\eta(t)\eta(t^{\prime})\rangle\propto T\delta_{ij}\delta(t-t^{\prime})$. Here, $T$ plays the role of an effective temperature that quantifies the strength of fluctuations in the system.
It is known that this equation of motion drives the system to a Gibbs distribution [8, 53, 54, 55]. Specifically, after a sufficiently long period of training, the probability distribution for a particular synaptic configuration ${W,b}$ is given by:
@@
where $\beta=\frac{1}{T}$, and the temperature $T$ arises from the strength of the stochastic noise in the training dynamics, as described above. In this paper, we adopt a framework wherein the synaptic weights of the DNN follow a Gibbs distribution, with the cost function serving as the energy. The partition function, denoted by $Z$, is defined as:
@@
where $\{W,b\}$ represents all possible configurations of synaptic weights and biases. [^4] Considering all configurations of the system implies accounting for an infinite amount of time in the training process. However, in most examples from statistical physics, it is widely observed that a sufficiently long experimental run ensures that the system follows the Gibbs distribution. Therefore, if the training time is sufficiently long, we expect our argument to apply to such a system.
@@ III Dynamic neurons
To address this difficulty, we promote the neurons $h_{i}^{(m)[l]}$ to additional dynamic degrees of freedom and ensure that Eq. (1) is dynamically satisfied. Instead of using iterated activation functions, we introduce a series of Dirac delta functions and simplify the exponent to the absolute square as follows:
$\begin{split}Z=\int dWdbdh\prod_{i,l}\bigg{[}&\exp\Big{[}-\beta\left(y^{[l]}_{i}-h^{(M)[l]}_{i}\right)^{2}\Big{]}\\ &\hskip 5.69054pt\prod_{m}\delta\left(h_{i}^{(m+1)[l]}-\sigma_{i}^{(m)[l]}\right)\bigg{]}.\end{split}$ (11)
Here, $i$ runs over the widths, $l$ runs over the data numbers, and $m$ runs over the layers. We refer to these new dynamic degrees of freedom, $h^{(m)[l]}$, as “dynamic neurons.” While this partition function appears straightforward, it does not yet simplify the computation.
@@
$\begin{split}Z=\int d&Wdbdh\lim_{v\to 0^{+}}\left(\frac{1}{v\sqrt{2\pi}}\right)^{n}\\ &\exp\Big{[}-\beta\sum_{i,l}\left(y_{i}^{[l]}-\sigma^{(M)[l]}_{i}\right)^{2}\\ &\qquad\quad-\beta_{v}\sum_{i,l,m}\left(h^{(m+1)[l]}_{i}-\sigma_{i}^{(m)[l]}\right)^{2}\Big{]},\end{split}$ (13)
@@
We observe that incorporating the constraint terms reveals the internal structure of a DNN within the Hamiltonian framework. By defining the new Hamiltonian [^7] If the cost function is defined using an $L^{p}$-norm, we can define the new Hamiltonian as $H_{\text{new}}=\sum\lambda|h-\sigma|^{p}$. as
@@
with $\lambda_{m}=\beta_{v}/\beta$ except for $\lambda_{M}=1$, we can denote Eq. (13) in the typical way of describing the partition function,
@@
where the limit over $v$ is omitted for simplicity. The form of this new Hamiltonian is the summation of interactions between the nearest layers [$m$-th layer and $(m+1)$-th layer]. In other words, interactions in the new Hamiltonian are restricted to the nearest degrees of freedom following the same pattern, unlike the original Hamiltonian, thereby implying locality.
@@
Although we introduced the infinite $\beta_{v}$ in the previous discussion, we will now study a more general system with an arbitrary positive $\beta_{v}$. When $\beta_{v}$ is sufficiently large, the system is expected to behave similarly to the infinite $\beta_{v}$ case. In other words, we are considering a generalization of deep learning in which the limit of large $\beta_{v}$ corresponds to the typical DNN.
@@
The emergence of this type of translation symmetry can simplify calculations and lead to interesting consequences. For example, in the one-dimensional and two-dimensional Ising models, the transfer matrix can be calculated explicitly due to translational symmetry. Therefore, it is expected that the new Hamiltonian can be analyzed by utilizing this symmetry. As an example, we perform the RG transformation in Sec. IV.
@@
Moreover, this system can be effectively described using principles from statistical physics, which are often employed to uncover the macroscopic behavior of complex systems composed of similarly interacting units.
With certain simplifying assumptions, this approach yields a Hamiltonian that closely resembles the spin glass model, a well-established framework in statistical physics. [^8] Several approaches have already introduced spin glass models to analyze DNNs under various assumptions [11, 12, 10, 13, 14, 15, 16, 17]. (See Appendix B.) This approach offers a new method for applying the spin glass model to gradient descent-based training in DNNs. Also, since we represented the training process of a general DNN using tools from statistical physics, we can inversely implement a system governed by such a Hamiltonian through the training of a DNN. (See Appendix A.)
@@ IV Renormalization Group Transformation
Since the new Hamiltonian resembles well-known statistical models, we can apply statistical physics tools to analyze the training of DNNs. A promising method in statistical physics is the RG transformation, which is widely used in fields such as condensed matter and high-energy physics. This technique systematically simplifies a system by progressively integrating out smaller-scale fluctuations, leading to a renormalized system with fewer degrees of freedom.
@@
Figure 3: RG transformation applied to the DNN using the DNA, resulting in reduced network depth. A notable similarity with the Ising model is apparent.
(i) The system consists of one neuron per layer ($N=1$) with a large depth ($M$ is large), and the training dataset contains a single data point ($L=1$). This assumption may be oversimplification and the example could be quite far from actual DNNs. In fact, the Modified National Institute of Standards and Technology (MNIST) [62] data consist of 60000 training examples ($L=60000$), and the universal approximation theorem prefers a sizable width with $N>1$ [63, 64, 65, 66, 67]. A lack of data can lead to insufficient training and overfitting, while insufficient depth may undermine the validity of the universal approximation theorem. It should be highlighted that our approach can be generally applied to DNNs with arbitrary $N,\ M,\ L$. Nevertheless, we make this assumption because our primary interest here lies in analytically demonstrating the statistical physics properties.
@@
(iii) The weights $W$ and biases $b$ are restricted to values of $-1$ or $1$, and the neurons $h$ are limited to values of 00 or $1$. While, in theory, $h$ should be integrated over all possible values, we expect the Gaussian-like behavior of the partition function to constrain $h$ to values near 00 or $1$. Although in general DNNs $W$ and $b$ can take arbitrary real numbers, we restrict their values here for simplicity. Interestingly, recent research has explored neural networks where parameters are similarly restricted to discrete values, such as $-1,1$ or $-1,0,1$. Despite these limitations, that model appears to achieve good performance [69, 70].
@@
As discussed in Sec. III, the Hamiltonian exhibits translational symmetry similar to that of the Ising model. Consequently, we can apply the RG transformation in a manner similar to its application in the Ising model. One approach involves summing over $W^{(m)}$, $b^{(m)}$, and $h^{(m)}$ for even-numbered layers in the partition function. After performing this summation, the partition function describes a system with fewer degrees of freedom. This real-space RG transformation method is known as decimation [71, 72], and the translational invariance or periodicity plays a crucial role in performing the decimation. [^10] Decimation is generally avoided for systems with dimensions higher than one because it fails to account for the renormalization of the magnitude of the degrees of freedom, preventing it from achieving a true fixed point in higher dimensions.
@@
The coupling constant $A$ has two critical aspects worth emphasizing. First, it introduces the interaction between neurons across different layers, effectively encoding the mechanism of information propagation within the system. Consequently, any changes in $A$ under RG transformation are expected to significantly influence the learning dynamics and overall behavior of the deep learning model. Second, this coupling term bears resemblance to the spin-spin interaction term in the Ising model. In the Ising model, the spin-spin interaction is the primary driver behind phase transitions, reflecting collective behavior changes at critical points. Similarly, in the context of the dynamic neuron system, $A$ is likely to play a pivotal role in phase transitions, potentially governing shifts in the system’s operational regimes or learning efficiency. Understanding this analogy provides a deeper insight into how critical phenomena might emerge in such neural systems.
Figure 4: RG flow near the fixed points $e^{A^{*}}=1$, demonstrating that RG flows converge to the plane composed of fixed points.
@@
Let us discuss the physical meaning of the RG transformation in the DNA. As discussed earlier, the introduction of dynamic neurons results in additional terms, $\lambda_{m}H_{m}$, which reflect the constraints,
@@
can be interpreted as new constraints between the nearest degrees of freedom in the renormalized system. As the depth of the DNN decreases by applying the RG transformation, new constraints for the system with reduced depth emerge. Therefore, the RG transformation suggests statistically equivalent constraints for the system with reduced depth. Theoretically, by applying these new constraints to the partition function, the introduced dynamic neurons can be integrated out. Subsequently, a renormalized DNN system without dynamic neurons can be obtained. In general, after applying the RG transformation, the constraints often become more complex, leading to multiple solutions or analytically unsolvable equations. However, even if we cannot solve them explicitly, identifying these constraints can aid in understanding the structure of DNNs. In this sense, the DNA is a valuable tool for comprehending the nature of DNNs. This argument can be generalized without relying on the simplifying assumptions (i)–(iii).
The discussions in this section are schematically summarized in Fig. 6. Because the RG transformation is a powerful technique in statistical physics, we aim to use it to analyze DNNs. However, directly applying the RG transformation to DNNs is challenging, so we employ the DNA. Due to the translational symmetry that emerges from the DNA, the RG transformation becomes much easier to apply. When the DNN undergoes an RG transformation using the DNA, it suggests new constraints for the system with reduced depth. By applying these constraints, one can obtain a renormalized DNN system.
DNNDNARenormalizedDNNRenormalizedDNARGRGIntroduceDynamicNeuronsApplyRenormalizedConstraints
Figure 6: This diagram illustrates how to develop a renormalized DNN using the DNA. Although not a rigorous explanation, it effectively conveys the essential concept. By introducing dynamic neurons, we obtain a new description of the DNN, and from the advantages of this new description, we derive a renormalized DNN with DNA. Finally, by applying renormalized constraints, the dynamic neurons can be integrated out, resulting in a renormalized DNN.
Before concluding this section, let us discuss how to relax the assumptions we introduced earlier in this section for the sake of demonstration. Using the decimation-based RG transformation may not be suitable when relaxing these assumptions. If assumption (i) is relaxed by increasing the sizes of $L$ and $N$, decimation can still be applied, but the number of couplings increases, making it challenging to analytically determine the fixed points. If assumption (ii) is generalized to allow different activation functions, assumption (iii) concerning the available values of the dynamic neuron must also be extended accordingly. If assumption (iii) is relaxed, expanding the available values of synaptic weights, biases, and dynamic neurons could result in an infinite number of possible values. In this case, polynomial expansion in decimation would no longer be feasible.
Nevertheless, these challenges arise from the difficulty of applying decimation, not from the limitations of the DNA itself. DNA can be applied regardless of assumptions (i), (ii), and (iii). Referring to the equations we presented in the previous section, extending the DNA to more general scenarios is straightforward. Statistical physics offers numerous RG transformation methods that do not rely on decimation. Applying these methods could enable further research on DNNs based on the DNA without these assumptions.